# Improved Approximation Algorithms for Minimizing the Total Weighted Completion Time of Coflows

### 20 Dec 2023 v2

@@ 4 Randomized Approximation Algorithm for the Coflow Scheduling Problem
$T=\max_{k\in\mathcal{K}}r_{k}+\max\left\{\max_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}L_{ik}+\max_{j\in\mathcal{J}}\sum_{k\in\mathcal{K}}L_{jk}\right\}-1.$
@@ 5 Deterministic Approximation Algorithm for the Coflow Scheduling Problem
$+\sum_{(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{F}_{i}\setminus\left\{(i,j,k)\right\}}\sum_{\ell=r_{k^{\prime}}}^{t-1}y_{i^{\prime}j^{\prime}k^{\prime}\ell}+\sum_{(i,j^{\prime},k^{\prime})<(i,j,k)}y_{ij^{\prime}k^{\prime}t}$
$+\sum_{(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{F}_{j}\setminus\left\{(i,j,k)\right\}}\sum_{\ell=r_{k^{\prime}}}^{t-1}y_{i^{\prime}j^{\prime}k^{\prime}\ell}+\sum_{(i^{\prime},j,k^{\prime})<(i,j,k)}y_{i^{\prime}jk^{\prime}t}$
@@
$+\sum_{(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{P}_{i}\cup\mathcal{P}_{j}}\sum_{\ell=r_{k^{\prime}}}^{t-1}x_{i^{\prime}j^{\prime}k^{\prime}\ell}d_{i^{\prime}j^{\prime}k^{\prime}}+\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{P}_{i}\cup\mathcal{P}_{j}\\ (i^{\prime},j^{\prime},k^{\prime})<(i,j,k)\end{subarray}}x_{i^{\prime}j^{\prime}k^{\prime}t}d_{i^{\prime}j^{\prime}k^{\prime}}$
$+\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},k^{\prime})\in\\ (\mathcal{F}_{i}\cup\mathcal{F}_{j})\setminus\left\{\mathcal{P}_{i}\cup\mathcal{P}_{j}\cup(i,j,k)\right\}\end{subarray}}\sum_{\ell=r_{k^{\prime}}}^{t-1}y_{i^{\prime}j^{\prime}k^{\prime}\ell}+\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},k^{\prime})\in\\ (\mathcal{F}_{i}\cup\mathcal{F}_{j})\setminus\left\{\mathcal{P}_{i}\cup\mathcal{P}_{j}\cup(i,j,k)\right\}\\ (i^{\prime},j^{\prime},k^{\prime})<(i,j,k)\end{subarray}}y_{i^{\prime}j^{\prime}k^{\prime}t}$
@@
$+\sum_{(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{P}_{i}\cup\mathcal{P}_{j}}\sum_{\ell=r_{k^{\prime}}}^{t-1}x_{i^{\prime}j^{\prime}k^{\prime}\ell}d_{i^{\prime}j^{\prime}k^{\prime}}+\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{P}_{i}\cup\mathcal{P}_{j}\\ (i^{\prime},j^{\prime},k^{\prime})<(i,j,k)\end{subarray}}x_{i^{\prime}j^{\prime}k^{\prime}t}d_{i^{\prime}j^{\prime}k^{\prime}}$
$+\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},k^{\prime})\in\\ (\mathcal{F}_{i}\cup\mathcal{F}_{j})\setminus\left\{\mathcal{P}_{i}\cup\mathcal{P}_{j}\right\}\end{subarray}}\sum_{\ell=r_{k^{\prime}}}^{t-1}y_{i^{\prime}j^{\prime}k^{\prime}\ell}+\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},k^{\prime})\in\\ (\mathcal{F}_{i}\cup\mathcal{F}_{j})\setminus\left\{\mathcal{P}_{i}\cup\mathcal{P}_{j}\right\}\\ (i^{\prime},j^{\prime},k^{\prime})<(i,j,k)\end{subarray}}y_{i^{\prime}j^{\prime}k^{\prime}t}$
be expectation completion time of flow $(i,j,k)$ to which flow $(i,j,k)$ has been assigned by $t$ when $(i,j,k)\in\mathcal{P}_{i}\cup\mathcal{P}_{j}$. The expected completion time of coflow $k$ is the maximum expected completion time among its flows. We have
$E_{\mathcal{P}_{i},\mathcal{P}_{j},x}[C_{k}]$ $=$ $\max\left\{A,B\right\}$ (9)
@@
$A$ $=$ $\max_{(i,j,k)\in\mathcal{F}_{k}\setminus\left(\mathcal{P}_{i}\cup\mathcal{P}_{j}\right)}\left\{\sum_{t=r_{k}}^{T}\frac{y_{ijkt}}{d_{ijk}}D(i,j,k,t)\right\},$
$B$ $=$ $\max_{(i,j,k)\in\mathcal{F}_{k}\cap\left(\mathcal{P}_{i}\cup\mathcal{P}_{j}\right)}\left\{E(i,j,k,t_{ijk})\right\}.$ (10)
@@ 6 Interval-indexed Linear Programming Relaxations
$\leq$ $d_{ijk}+\sum_{(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{P}_{i}\setminus\left\{(i,j,k)\right\}}\frac{1}{2}y_{i^{\prime}j^{\prime}k^{\prime}0}+\sum_{(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{P}_{j}\setminus\left\{(i,j,k)\right\}}\frac{1}{2}y_{i^{\prime}j^{\prime}k^{\prime}0}$
@@
$C(i,j,k,0)$ $=$ $d_{ijk}+\sum_{(i,j^{\prime},k^{\prime})<(i,j,k)}y_{ij^{\prime}k^{\prime}0}+\sum_{(i^{\prime},j,k^{\prime})<(i,j,k)}y_{i^{\prime}jk^{\prime}0}$
@@
$E[C_{k}]$ $=$ $\max_{(i,j,k)\in\mathcal{F}_{k}}\left\{\sum_{\ell=r_{k}}^{T}\frac{y_{ijk\ell}\cdot|I_{\ell}|}{d_{ijk}}C(i,j,k,\ell)\right\}.$
@@
$D(i,j,k,0)$ $=$ $d_{ijk}+\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{P}_{i}\cup\mathcal{P}_{j}\\ (i^{\prime},j^{\prime},k^{\prime})<(i,j,k)\end{subarray}}x_{i^{\prime}j^{\prime}k^{\prime}0}d_{i^{\prime}j^{\prime}k^{\prime}}+\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},k^{\prime})\in\\ (\mathcal{F}_{i}\cup\mathcal{F}_{j})\setminus\left\{\mathcal{P}_{i}\cup\mathcal{P}_{j}\cup(i,j,k)\right\}\\ (i^{\prime},j^{\prime},k^{\prime})<(i,j,k)\end{subarray}}y_{i^{\prime}j^{\prime}k^{\prime}0}$
@@
$E(i,j,k,0)$ $=$ $d_{ijk}+\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{P}_{i}\cup\mathcal{P}_{j}\\ (i^{\prime},j^{\prime},k^{\prime})<(i,j,k)\end{subarray}}x_{i^{\prime}j^{\prime}k^{\prime}0}d_{i^{\prime}j^{\prime}k^{\prime}}+\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},k^{\prime})\in\\ (\mathcal{F}_{i}\cup\mathcal{F}_{j})\setminus\left\{\mathcal{P}_{i}\cup\mathcal{P}_{j}\right\}\\ (i^{\prime},j^{\prime},k^{\prime})<(i,j,k)\end{subarray}}y_{i^{\prime}j^{\prime}k^{\prime}0}\cdot|I_{0}|$
@@
The expected completion time of coflow $k$ is the maximum expected completion time among its flows. We have
$E_{\mathcal{P}_{i},\mathcal{P}_{j},x}[C_{k}]$ $=$ $\max\left\{A,B\right\}$
@@
$A$ $=$ $\max_{(i,j,k)\in\mathcal{F}_{k}\setminus\left(\mathcal{P}_{i}\cup\mathcal{P}_{j}\right)}\left\{\sum_{\begin{subarray}{c}\ell=0\\ (1+\eta)^{\ell-1}\geq r_{k}\end{subarray}}^{T}\frac{y_{ijk\ell}\cdot|I_{\ell}|}{d_{ijk}}D(i,j,k,\ell)\right\},$
$B$ $=$ $\max_{(i,j,k)\in\mathcal{F}_{k}\cap\left(\mathcal{P}_{i}\cup\mathcal{P}_{j}\right)}\left\{E(i,j,k,\ell_{ijk})\right\}.$ (32)
@@
$\leq$ $d_{ijk}+\sum_{(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{P}_{i}\setminus\left\{(i,j,k)\right\}}y_{i^{\prime}j^{\prime}k^{\prime}0}+\sum_{(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{P}_{j}\setminus\left\{(i,j,k)\right\}}y_{i^{\prime}j^{\prime}k^{\prime}0}$

### 24 Dec 2023 v3

@@ 4 Randomized Approximation Algorithm for the Coflow Scheduling Problem
$T=\max_{k\in\mathcal{K}}r_{k}+\max_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}L_{ik}+\max_{j\in\mathcal{J}}\sum_{k\in\mathcal{K}}L_{jk}-1.$
@@ 5 Deterministic Approximation Algorithm for the Coflow Scheduling Problem
$+\sum_{(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{F}_{i}\setminus\left\{(i,j,k)\right\}}\sum_{\ell=r_{k^{\prime}}}^{t-1}y_{i^{\prime}j^{\prime}k^{\prime}\ell}$
$+\sum_{(i,j^{\prime},k^{\prime})<(i,j,k)}y_{ij^{\prime}k^{\prime}t}$
$+\sum_{(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{F}_{j}\setminus\left\{(i,j,k)\right\}}\sum_{\ell=r_{k^{\prime}}}^{t-1}y_{i^{\prime}j^{\prime}k^{\prime}\ell}$
$+\sum_{(i^{\prime},j,k^{\prime})<(i,j,k)}y_{i^{\prime}jk^{\prime}t}$
@@
$+\sum_{(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{P}_{i}\cup\mathcal{P}_{j}}\sum_{\ell=r_{k^{\prime}}}^{t-1}x_{i^{\prime}j^{\prime}k^{\prime}\ell}d_{i^{\prime}j^{\prime}k^{\prime}}$
$+\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{P}_{i}\cup\mathcal{P}_{j}\\ (i^{\prime},j^{\prime},k^{\prime})<(i,j,k)\end{subarray}}x_{i^{\prime}j^{\prime}k^{\prime}t}d_{i^{\prime}j^{\prime}k^{\prime}}$
$+\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},k^{\prime})\in\\ (\mathcal{F}_{i}\cup\mathcal{F}_{j})\setminus\left\{\mathcal{P}_{i}\cup\mathcal{P}_{j}\cup(i,j,k)\right\}\end{subarray}}\sum_{\ell=r_{k^{\prime}}}^{t-1}y_{i^{\prime}j^{\prime}k^{\prime}\ell}$
$+\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},k^{\prime})\in\\ (\mathcal{F}_{i}\cup\mathcal{F}_{j})\setminus\left\{\mathcal{P}_{i}\cup\mathcal{P}_{j}\cup(i,j,k)\right\}\\ (i^{\prime},j^{\prime},k^{\prime})<(i,j,k)\end{subarray}}y_{i^{\prime}j^{\prime}k^{\prime}t}$
@@
$+\sum_{(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{P}_{i}\cup\mathcal{P}_{j}}\sum_{\ell=r_{k^{\prime}}}^{t-1}x_{i^{\prime}j^{\prime}k^{\prime}\ell}d_{i^{\prime}j^{\prime}k^{\prime}}$
$+\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{P}_{i}\cup\mathcal{P}_{j}\\ (i^{\prime},j^{\prime},k^{\prime})<(i,j,k)\end{subarray}}x_{i^{\prime}j^{\prime}k^{\prime}t}d_{i^{\prime}j^{\prime}k^{\prime}}$
$+\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},k^{\prime})\in\\ (\mathcal{F}_{i}\cup\mathcal{F}_{j})\setminus\left\{\mathcal{P}_{i}\cup\mathcal{P}_{j}\right\}\end{subarray}}\sum_{\ell=r_{k^{\prime}}}^{t-1}y_{i^{\prime}j^{\prime}k^{\prime}\ell}$
$+\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},k^{\prime})\in\\ (\mathcal{F}_{i}\cup\mathcal{F}_{j})\setminus\left\{\mathcal{P}_{i}\cup\mathcal{P}_{j}\right\}\\ (i^{\prime},j^{\prime},k^{\prime})<(i,j,k)\end{subarray}}y_{i^{\prime}j^{\prime}k^{\prime}t}$
be expectation completion time of flow $(i,j,k)$ to which flow $(i,j,k)$ has been assigned by $t$ when $(i,j,k)\in\mathcal{P}_{i}\cup\mathcal{P}_{j}$. Let $\mathcal{P}\subseteq F$ represent subsets of flows that have already been assigned the switch-interval pair. The expected completion time of coflow $k$ is the maximum expected completion time among its flows. We have
$E_{\mathcal{P},x}[C_{k}]$ $=$ $\max\left\{A,B\right\}$ (11)
@@
$A$ $=$ $\max_{(i,j,k)\in\mathcal{F}_{k}\setminus\mathcal{P}}\left\{\sum_{t=r_{k}}^{T}\frac{y_{ijkt}}{d_{ijk}}D(i,j,k,t)\right\},$
$B$ $=$ $\max_{(i,j,k)\in\mathcal{F}_{k}\cap\mathcal{P}}\left\{E(i,j,k,t_{ijk})\right\}.$ (12)
@@ 6 Interval-indexed Linear Programming Relaxations
$\leq$ $d_{ijk}+\sum_{(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{P}_{i}\setminus\left\{(i,j,k)\right\}}\frac{1}{2}y_{i^{\prime}j^{\prime}k^{\prime}0}$
$+\sum_{(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{P}_{j}\setminus\left\{(i,j,k)\right\}}\frac{1}{2}y_{i^{\prime}j^{\prime}k^{\prime}0}$
@@
$C(i,j,k,0)$ $=$ $d_{ijk}+\sum_{(i,j^{\prime},k^{\prime})<(i,j,k)}y_{ij^{\prime}k^{\prime}0}$
$+\sum_{(i^{\prime},j,k^{\prime})<(i,j,k)}y_{i^{\prime}jk^{\prime}0}$
@@
$E[C_{k}]$ $=$ $\max_{(i,j,k)\in\mathcal{F}_{k}}\left\{\sum_{\begin{subarray}{c}\ell=0\\ (1+\eta)^{\ell-1}\geq r_{k}\end{subarray}}^{L}\frac{y_{ijk\ell}\cdot|I_{\ell}|}{d_{ijk}}C(i,j,k,\ell)\right\}.$
@@
$D(i,j,k,0)$ $=$ $d_{ijk}+\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{P}_{i}\cup\mathcal{P}_{j}\\ (i^{\prime},j^{\prime},k^{\prime})<(i,j,k)\end{subarray}}x_{i^{\prime}j^{\prime}k^{\prime}0}d_{i^{\prime}j^{\prime}k^{\prime}}$
$+\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},k^{\prime})\in\\ (\mathcal{F}_{i}\cup\mathcal{F}_{j})\setminus\left\{\mathcal{P}_{i}\cup\mathcal{P}_{j}\cup(i,j,k)\right\}\\ (i^{\prime},j^{\prime},k^{\prime})<(i,j,k)\end{subarray}}y_{i^{\prime}j^{\prime}k^{\prime}0}$
@@
$E(i,j,k,0)$ $=$ $d_{ijk}+\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{P}_{i}\cup\mathcal{P}_{j}\\ (i^{\prime},j^{\prime},k^{\prime})<(i,j,k)\end{subarray}}x_{i^{\prime}j^{\prime}k^{\prime}0}d_{i^{\prime}j^{\prime}k^{\prime}}$
$+\sum_{\begin{subarray}{c}(i^{\prime},j^{\prime},k^{\prime})\in\\ (\mathcal{F}_{i}\cup\mathcal{F}_{j})\setminus\left\{\mathcal{P}_{i}\cup\mathcal{P}_{j}\right\}\\ (i^{\prime},j^{\prime},k^{\prime})<(i,j,k)\end{subarray}}y_{i^{\prime}j^{\prime}k^{\prime}0}\cdot|I_{0}|$
@@
Let $\mathcal{P}\subseteq F$ represent subsets of flows that have already been assigned the switch-interval pair. The expected completion time of coflow $k$ is the maximum expected completion time among its flows. We have
$E_{\mathcal{P},x}[C_{k}]$ $=$ $\max\left\{A,B\right\}$
@@
$A$ $=$ $\max_{(i,j,k)\in\mathcal{F}_{k}\setminus\mathcal{P}}\left\{\sum_{\begin{subarray}{c}\ell=0\\ (1+\eta)^{\ell-1}\geq r_{k}\end{subarray}}^{L}\frac{y_{ijk\ell}\cdot|I_{\ell}|}{d_{ijk}}D(i,j,k,\ell)\right\},$
$B$ $=$ $\max_{(i,j,k)\in\mathcal{F}_{k}\cap\mathcal{P}}\left\{E(i,j,k,\ell_{ijk})\right\}.$ (37)
@@
$\leq$ $d_{ijk}+\sum_{(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{P}_{i}\setminus\left\{(i,j,k)\right\}}y_{i^{\prime}j^{\prime}k^{\prime}0}$
$+\sum_{(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{P}_{j}\setminus\left\{(i,j,k)\right\}}y_{i^{\prime}j^{\prime}k^{\prime}0}$